OpenStudy (anonymous):

Convert the equation to the standard form for a parabola by completing the square on x or y as appropriate. y2 + 2y - 2x - 3 = 0

OpenStudy (anonymous):

A. (y + 1)2 = 2(x + 2) B. (y - 1)2 = -2(x + 2) C. (y + 1)2 = 2(x - 2) D. (y - 1)2 = 2(x + 2)